OpenStudy (anonymous):

PLEASE HELP! SUPER IMPORTANT! The 1st, 5th, and 13th terms of an arithmetic sequence are the first three terms of a geometric sequence with a common ratio of 2. If the 21st term of the arithmetic sequence is 72, calculate the sum of the first 10 terms of the geometric sequence.

OpenStudy (princeharryyy):

let the terms of gp be a,ar,ar^2,r=2 given, term21 = 72 => a+(21-1)d=72 gives a+ 20d =72, term5=ar given => a+(5-1)d=2a given in question, =>a+4d=2a gives a=4d, put the value of d in term21 eqution which gives a =12, use gp formula as now u will kno the terms of gp, ans gives (12/11)*((12^10)-1)

OpenStudy (anonymous):

thank you!

OpenStudy (princeharryyy):

welcome :)